OpenStudy (anonymous):

solve( -18x^3+17x+6)/(3x+2)

OpenStudy (anonymous):

HELP

OpenStudy (opcode):

Is this multiple choice?

OpenStudy (anonymous):

try change (-18x^3+17x+6) into (3x+2)(ax+b) form

OpenStudy (wolfe8):

How can you solve this? There is no equation.

OpenStudy (opcode):

I believe she wants the quotient and remainder.

OpenStudy (anonymous):

@wolfe8 in this case solve would mean simplify

OpenStudy (wolfe8):

Simplify the following: (-18 x^3+17 x+6)/(3 x+2) Factor a minus sign out of -18 x^3+17 x+6. Factor -1 out of -18 x^3+17 x+6: -(18 x^3-17 x-6)/(3 x+2) Find all linear factors of 18 x^3-17 x-6 via the rational root theorem. Do this by finding rational roots. The candidates are x = ±p/q for all p that are divisors of the constant term -6 and for all q that are divisors of the leading coefficient 18. The possible rational roots of 18 x^3-17 x-6 are x = ±1/18, x = ±1/9, x = ±2/9, x = ±1/6, x = ±1/3, x = ±2/3, x = ±1/2, x = ±3/2, x = ±1, x = ±2, x = ±3, x = ±6. Of these, x = -2/3 is a root. This gives 3 x+2 as all linear factors: -(((3 x+2) (18 x^3-17 x-6))/(3 x+2))/(3 x+2) Divide 3 x+2 into 18 x^3-17 x-6. | | 3 x | + | 2 | | 6 x^2 | - | 4 x | - | 3 18 x^3 | + | 0 x^2 | - | 17 x | - | 6 18 x^3 | + | 12 x^2 | | | | | | -12 x^2 | - | 17 x | | | | -12 x^2 | - | 8 x | | | | | | -9 x | - | 6 | | | | -9 x | - | 6 | | | | | | 0: -(6 x^2-4 x-3 (3 x+2))/(3 x+2) Cancel common terms in the numerator and denominator of -((3 x+2) (6 x^2-4 x-3))/(3 x+2). -((3 x+2) (6 x^2-4 x-3))/(3 x+2) = (3 x+2)/(3 x+2)×-(6 x^2-4 x-3) = -(6 x^2-4 x-3): Answer: | | -(6 x^2-4 x-3)

OpenStudy (opcode):

Much neater:

OpenStudy (opcode):

\[9x + 6 - 9x + 6 = 0\] So the remainder is zero, the and the quotient is: \[-6x^2 + 4x +3\]